OpenStudy (anonymous):

Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers.

OpenStudy (anonymous):

x+2=x+2

OpenStudy (anonymous):

Can it be like that?

OpenStudy (anonymous):

you must prove it, so subtract 2 from both sides.

OpenStudy (anonymous):

SO, x+2=x+2 x+2-2=x+2-2 x=x {which is TRUE for all real numbers}

OpenStudy (anonymous):

correct